# The SPHEREx Satellite Mission

### v1

@@ Abstract
SPHEREx, a NASA explorer satellite launched on 11 March 2025, is carrying out the first all-sky near-infrared spectral survey. The satellite observes in 102 spectral bands from 0.75 to 5.0 with a resolving power ranging from in 62 pixels. The observatory obtains a depth of 19.5 - 19.9 AB mag for and 17.8 - 18.8 AB mag for after mapping the full sky four times over two years. Scientifically, SPHEREx will produce a large galaxy redshift survey over the full sky, intended to constrain the amplitude of inflationary non-Gaussianity. The observations will produce two deep spectral maps near the ecliptic poles that will use intensity mapping to probe the evolution of galaxies over cosmic history. By mapping the depth of infrared absorption features over the Galactic plane, SPHEREx will comprehensively survey the abundance and composition of water and other biogenic ice species in the interstellar medium. The initial data are rapidly released in the form of spectral images to the public. The project will release specialized data products over the life of the mission as the surveys proceed. The science team will also produce specialized spectral catalogs on planet-bearing and low-mass stars, solar system objects, and galaxy clusters 3 years after launch. We describe the design of the instrument and spacecraft, which flow from the core science requirements. Finally, we present an initial evaluation of the satellite’s in-flight performance and key characteristics.
@@ 1 Introduction
The Spectro-Photometer for the History of the universe, Epoch of Reionization, and Ices Explorer (SPHEREx) is designed to map the entire sky at near-infrared wavelengths () in low-resolution spectroscopy () with moderate 6″ angular resolution. SPHEREx follows a succession of infrared all-sky survey programs, starting with the InfraRed Astronomical Satellite (IRAS) (Neugebauer et al., 1984), the Diffuse InfraRed Background Experiment (DIRBE) on the COBE satellite (Silverberg et al., 1993), 2MASS (Skrutskie et al., 2006), AKARI (Murakami et al., 2007), and the Wide-field Infrared Survey Explorer (WISE) (Wright et al., 2010), all of which mapped the full sky in broad photometric bands. SPHEREx combines large-format sensitive infrared detector arrays, an entirely passive cooling design, a wide-field off-axis telescope, and the novel use of Linear Variable Filter (LVF) spectrometers to provide the first all-sky near-infrared spectroscopic survey. SPHEREx will produce 4 complete all-sky spectral maps over its 2-year baseline mission.
@@ 2.1 Inflationary Cosmology
SPHEREx tests the inflationary paradigm in modern cosmology, which was proposed to explain the large-scale homogeneity and geometric flatness of the universe (Guth, 1981; Linde, 1982; Albrecht & Steinhardt, 1982; Starobinsky, 1980). The basic predictions of inflation have been verified in increasingly precise measurements (Planck Collaboration et al., 2020a), specifically a near-flat spatial geometry on large scales with initial density perturbations that are nearly Gaussian, phase-synchronous, and nearly, but not precisely, scale-invariant. The physical process driving inflation, however, remains elusive.
@@
Observations of non-Gaussianity in the primordial fluctuations, quantified by the parameters, describe the departure from a Gaussian bell curve (Komatsu et al., 2005), and provide a unique test between multi-field and single-field models. Inflation models driven by multiple fields generally predict a local non-Gaussianity while simpler models with a single field generally predict (Alvarez et al., 2014; de Putter et al., 2017) (See Figure 1). For simplicity, we will write for from now on.
CMB-based measurements currently constrain (Planck Collaboration et al., 2020b), close to the limits set by cosmic variance (Baumann et al., 2009). Because large-scale structure galaxy surveys map the universe in 3 dimensions, they access many more modes than a 2-dimensional CMB map, and can thus probe non-Gaussianity to much lower levels than are possible with CMB measurements. Current constraints from galaxy surveys are now starting to approach the accuracy of the CMB (Chaussidon et al., 2025).
SPHEREx probes using the power spectrum and bispectrum, the Fourier transforms of the 2-point and 3-point galaxy correlation function, respectively. The relative sensitivity of these two measures is similar. Combined, they provide a powerful internal consistency test, because the two statistical measures are quite different. For the power spectrum, the effect of is most prominent on large spatial scales because of the scale dependent galaxy bias induced by (Dalal et al., 2008). Its accuracy is driven by maximizing the number of redshifts, which can have low accuracy , over a large cosmological volume. The bispectrum is driven by having multiple galaxy samples with smaller high-accuracy redshifts up to to measure squeezed triangles with legs having a large spatial wavenumber (Heinrich et al., 2024).
@@ 2.2 History of Galaxy Formation
SPHEREx probes the history of galaxy formation by studying the spatial anisotropy of the EBL, the aggregate glow of which contains emission from all sources integrated over redshift. By mapping structures from degree to 6″ angular scales, SPHEREx accesses signals from linear (2-halo) clustering, non-linear (1-halo) clustering, and Poisson (shot noise) fluctuations (e.g. Asgari et al. (2023), Kovetz et al. (2017), Cooray & Sheth (2002)). Linear clustering refers to scales between dark matter halos that can be traced back to primordial fluctuations, while non-linear clustering refers to smaller scales within a halo where the non-linear effects of gravity and star formation are important. Combined, these provide measures of the total light production, the brightness of clustered galaxies and intrahalo light (IHL) associated with stripped stars in dark matter halos, and the ensemble properties of galaxies above a selected brightness.
On large spatial scales, galaxies trace dark matter over-densities through linear clustering. The linear clustering signal is of particular scientific interest because it provides an integrated measure of luminosity , where is the bias of emitting sources and is the total light emitted at a specific redshift from all sources, including individually faint galaxies and diffuse components that may be missed in galaxy-counting surveys. SPHEREx has two ways to extract the redshift history of emission. Because each wave band is sensitive to a different range of redshifts, the cross-spectrum between two bands partially correlates and provides information on the history of light production. The numerous cross-spectra thus enable fitting a suite of parameters to describe various emission components and their luminosity history (Feng et al., 2019). An alternative approach is to cross-correlate SPHEREx maps with galaxy redshifts (Cheng & Chang, 2022), which provides an unambiguous redshift separation, but is limited to available redshift surveys.
@@ 2.3 Interstellar Ices
Key biogenic molecules, such as water (H2O), carbon dioxide (CO2), carbon monoxide (CO) and methanol (CH3OH), are locked in ices (Boogert et al., 2011, 2013; Öberg et al., 2011) on the mantles of interstellar dust grains, in amounts predicted in some cases to far exceed those in the gas phase (Hollenbach et al., 2009). Ices play a major role in forming planetesimals within disks, and are a key source of water and organic molecules for newly forming planets. Grain mantles are the sites of active surface chemistry (Tielens & Hagen, 1982; Cuppen et al., 2009), and thus one might expect ice composition to depend on their environment. The limited number of existing ice spectra is insufficient to determine whether the ices in protoplanetary disks, including our early Solar System, are mainly inherited from their progenitor clouds (Dhooghe et al., 2017), or whether ice composition evolves over time (Aikawa et al., 2012).
SPHEREx addresses these questions by conducting a comprehensive survey of ices to better understand their abundances, compositions, and how they evolve during the early phases of star and planet formation. We selected a list of more than 9 million targets based on 2MASS and WISE 3.4 and 4.6 µm catalogs using three criteria: 1) a line-of-sight extinction ; 2) a WISE 4.6 m flux brighter than 15 AB mag to ensure a high signal-to-noise ratio in each SPHEREx spectral channel; and 3) targets that are separated from any neighboring source contributing more than 1 % of their flux by at least 7″. The SPHEREx ices target list (Ashby et al., 2023) will greatly expand – by more than a factor of 10,000 – the database of ice absorption spectra previously obtained by ISO, AKARI, Spitzer and JWST, significantly advancing our understanding of how ices form and evolve as gas and dust transition from diffuse to dense interstellar clouds, to circumstellar envelopes, and ultimately to protoplanetary disks (Melnick et al. (2025); see Figure 3).
@@ 2.4 All-Sky Spectral Survey
All-sky surveys (e.g. IRAS, COBE, AKARI, WISE) play a major role in advancing the field of astronomy, producing versatile legacy archives that prove valuable for decades. SPHEREx extends this tradition by providing the first all-sky near-infrared spectral survey. SPHEREx’s rich spectral database supports numerous scientific investigations (Doré et al., 2016). These range from studies of main sequence, low-mass, and evolved stars, to spectral catalogs of galaxies, to surveys of solar system objects, to observations of clusters, to Galactic maps of line and PAH emission features. The all-sky survey complements recent and forthcoming astrophysics satellites (Doré et al., 2018), including JWST, TESS, Gaia, Euclid, eROSITA and Roman.
@@ 2.5 Legacy Catalogs
In addition to data products that support the 3 core science themes (see §2.1 - §2.3), the SPHEREx science team is curating 3 spectral catalogs that benefit from the team’s specialized knowledge of the instrument: asteroids and comets; planet-bearing and low-mass stars; and clusters of galaxies. The solar system objects catalog of asteroids and comets consists of objects with known, but changing positions. Ivezić et al. (2022) estimates that SPHEREx will deliver meaningful flux measurements for about 100,000 asteroids, of which 10,000 objects will have high-quality spectra. Because the spectral observations are distributed over time, care must be taken to account for the effects of rotation and changing distance. The wavelengths span from scattered sunlight to thermal emission and include characteristic features of olivine, pyroxene, hydroxyl, water ice, and organics, and contain information on thermal properties that complement broadband photometry with the upcoming NEO Surveyor satellite (Mainzer et al., 2023).
SPHEREx will provide spectra to supplement external photometry from Gaia, GALEX, 2MASS and WISE for over 600,000 main sequence stars surveyed for exoplanets by the transit missions Kepler, K2 and TESS. These stellar SEDs will help determine best-fit effective temperatures, surface gravities, and [Fe/H] ratios to percent or sub-percent precision. When combined with Gaia astrometry, we can derive bolometric luminosities, proper motions, stellar radii and extinctions. This combination of data significantly improves planetary radius estimates for many transiting planets, by reducing the error from uncertainty in the stellar radius, which can otherwise dominate, especially for cool stars (Stevens et al., 2018). SPHEREx will also compile a spectral atlas of 10,000 low-mass stars, from late M dwarfs and all accessible brown dwarfs down to the coolest Y dwarfs. Low-resolution near-infrared spectroscopy is crucial for studying these cool stellar objects, which can have rich spectra due to their complex atmospheres. The observations will use Gaia astrometry to account for significant proper motions.
@@ 3 Science to Requirements
SPHEREx is designed to fulfill the science requirements of its three primary scientific objectives. In short, the inflationary cosmology theme is driven by having sufficient high-accuracy and low-accuracy redshifts to constrain the non-Gaussianity parameter through the bispectrum (Heinrich et al., 2024) and power spectrum, respectively. These in turn set requirements for spectral resolving power and point source sensitivity in bands 1-4 (see Table 2). The history of galaxy formation theme can carry out its science with the same resolving power and sensitivity, but must control noise and systematic errors on degree angular scales in order to accurately map extended cosmological structures. The interstellar ices science theme studies bright Galactic sources, but requires higher resolving power, with wavelength coverage out to 5 to spectrally separate ice absorption features. We describe the flow of engineering requirements from science in abbreviated form, but note that extensive simulations were carried out to justify these choices during the mission development.
@@ 3.1 Spectral Coverage and Resolving Power
The four short-wavelength bands were designed for the cosmological redshift survey, with resolving powers of (see Table 2). We carried out simulations (Stickley et al., 2016) of SPHEREx observations on the COSMOS field, using realistic galaxy spectra and including the effects of spectral confusion from faint galaxies below SPHEREx’s sensitivity. These studies indicated that we could obtain redshifts with accuracy up to , required for the bispectrum analysis, using only continuum features. The most recent study of SPHEREx redshifts (Feder et al., 2024) confirms our choice of resolving power is appropriate, and indicates we could enhance the number of high-accuracy redshifts by incorporating emission lines. The galaxy formation theme can be accomplished with the wavelength and spectral resolution choices for cosmology.
@@ 3.2 Point Source Sensitivity
The inflationary cosmology theme constrains inflationary non-Gaussianity through measurements of the power spectrum and bispectrum with a large catalog of galaxies with a range of redshift accuracy (Doré et al., 2014). The number density of galaxies as a function of redshift, their clustering biases, and their redshift accuracy set the measurement error on . The forecasting methodology used for the power spectrum is described in Doré et al. (2014) and de Putter & Doré (2017). The forecasting methodology used for the bispectrum follows Heinrich et al. (2024) and includes an accurate handling of redshift errors, but excludes the non-Gaussian contribution to the covariance matrix (Biagetti et al., 2022; Salvalaggio et al., 2024) and the covariance between the power spectrum and bispectrum (de Putter, 2018). It also assumes the galaxy biases universality relation (Dalal et al., 2008; Barreira, 2020). All these effects, when included, might reduce the sensitivity by tens of percent. However, a full analysis incorporating a multi-tracer analysis with optimal samples may substantially improve the measurement error on from our initial forecast.
Feder et al. (2024) studied the number and accuracy of redshifts using a template fitting method, which strongly depends on point source sensitivity in bands 1-4. In order to achieve the accuracy needed to meaningfully discriminate between single-field and multi-field inflation models, we set a requirement that the point source sensitivity exceed 18.4 AB mag () per spectral channel (see §4.2) in bands 1-4 once observations are complete (see Figure 5).

### v2

@@ Abstract
SPHEREx, a NASA Explorer satellite launched on 11 March 2025, is carrying out the first all-sky near-infrared spectral survey. The satellite observes in 102 spectral bands from 0.75 to 5.0 with a resolving power ranging from in 62 pixels. The observatory obtains a depth of 19.5 - 19.9 AB mag for with and 17.8 - 18.8 AB mag for with after mapping the full sky four times over two years. Scientifically, SPHEREx will produce a large galaxy redshift survey over the full sky to constrain the amplitude of inflationary non-Gaussianity. The observations will produce two deep spectral maps near the ecliptic poles that use intensity mapping to probe the evolution of galaxies over cosmic history. By mapping the depth of infrared absorption features over the Galactic plane, SPHEREx will comprehensively survey the abundance and composition of water and other biogenic ice species in the interstellar medium. The project will release initial data rapidly in the form of spectral images, and specialized data products over the life of the mission as the surveys proceed. The science team will also produce spectral catalogs of planet-bearing and low-mass stars, solar system objects, and galaxy clusters 3 years after launch. We describe the design of the instrument and spacecraft, which flow from the core science requirements. Finally, we present an initial evaluation of the satellite’s in-flight performance and key characteristics.
@@
The Spectro-Photometer for the History of the universe, Epoch of Reionization, and Ices Explorer (SPHEREx) is designed to map the entire sky at near-infrared wavelengths () in low-resolution spectroscopy () with moderate 6″ angular resolution. SPHEREx (see Figure 1 follows a succession of infrared all-sky survey programs, starting with the InfraRed Astronomical Satellite (IRAS) (Neugebauer84), the Diffuse InfraRed Background Experiment (DIRBE) on the COBE satellite (Silverberg93), 2MASS (Skrutskie06), AKARI (Murakami07), and the Wide-field Infrared Survey Explorer (WISE) (Wright10), all of which mapped the full sky in broad photometric bands. SPHEREx combines large-format sensitive infrared detector arrays, an entirely passive cooling design, a wide-field off-axis telescope, and the novel use of Linear Variable Filter (LVF) spectrometers to provide the first all-sky near-infrared spectroscopic survey. SPHEREx will produce 4 complete all-sky spectral maps over its 2-year baseline mission.
@@ 2.1 Inflationary Cosmology
SPHEREx tests the inflationary paradigm in modern cosmology, which was proposed to explain the large-scale homogeneity and geometric flatness of the universe (guth81; linde82; albrecht82; starobinsky80). The basic predictions of inflation have been verified in increasingly precise measurements (planck_VI_20), specifically a near-flat spatial geometry on large scales with initial density perturbations that are nearly Gaussian, phase-synchronous, and nearly, but not precisely, scale-invariant. The physical process driving inflation, however, remains elusive.
@@
Observations of non-Gaussianity in the primordial fluctuations, quantified by the parameters, describe the departure from a Gaussian bell curve (komatsu05), and provide a unique test between multi-field and single-field models. Inflation models driven by multiple fields generally predict a local non-Gaussianity while simpler models with a single field generally predict (alvarez14; dePutter:2016trg) (See Figure 2). For simplicity, we will write for from now on.
CMB-based measurements currently constrain (planck_IX_20), close to the limits set by cosmic variance (baumann09). Because large-scale structure galaxy surveys map the universe in 3 dimensions, they access many more modes than a 2-dimensional CMB map, and can thus probe non-Gaussianity to much lower levels than are possible with CMB measurements. Current constraints from galaxy surveys are now starting to approach the accuracy of the CMB (Chaussidon_2025).
SPHEREx probes using the power spectrum and bispectrum, the Fourier transforms of the 2-point and 3-point galaxy correlation function, respectively. The relative sensitivity of these two measures is similar. Combined, they provide a powerful internal consistency test, because the two statistical measures are quite different. For the power spectrum, the effect of is most prominent on large spatial scales because of the scale dependent galaxy bias induced by (Dalal:2007cu). Its accuracy is driven by maximizing the number of redshifts, which can have low accuracy , over a large cosmological volume. The bispectrum is driven by having multiple galaxy samples with smaller high-accuracy redshifts up to to measure squeezed triangles with legs having a large spatial wavenumber (Heinrich24).
In addition to constraints on , we expect SPHEREx clustering measurements, with several hundred million galaxies over the full sky, to lead to new insights on the properties of dark energy, the curvature of the Universe, the mass of neutrinos (Dore14), and so-called general relativistic effects (Wen:2024).
@@ 2.2 History of Galaxy Formation
SPHEREx probes the history of galaxy formation by studying the spatial anisotropy of the EBL, the aggregate glow of which contains emission from all sources integrated over redshift. By mapping structures from degree to 6″ angular scales, SPHEREx accesses signals from linear (2-halo) clustering, non-linear (1-halo) clustering, and Poisson (shot noise) fluctuations (e.g. Asgari23, kovetz17, COORAY02). Linear clustering refers to scales between dark matter halos that can be traced back to primordial fluctuations, while non-linear clustering refers to smaller scales within a halo where the non-linear effects of gravity and star formation are important. Combined, these provide measures of the total light production, the brightness of clustered galaxies and intrahalo light (IHL) associated with stripped stars in dark matter halos, and the ensemble properties of galaxies above a selected brightness.
On large spatial scales, galaxies trace dark matter over-densities through linear clustering. The linear clustering signal is of particular scientific interest because it provides an integrated measure of luminosity , where is the bias of emitting sources and is the total light emitted at a specific redshift from all sources, including individually faint galaxies and diffuse components that may be missed in galaxy-counting surveys. SPHEREx has two ways to extract the redshift history of emission. Because each wave band is sensitive to a different range of redshifts, the cross-spectrum between two bands partially correlates and provides information on the history of light production. The numerous cross-spectra thus enable fitting a suite of parameters to describe various emission components and their luminosity history (feng19). An alternative approach is to cross-correlate SPHEREx maps with galaxy redshifts (cheng22), which provides an unambiguous redshift separation, but is limited to available redshift surveys.
@@ 2.3 Interstellar Ices
Key biogenic molecules, such as water (H2O), carbon dioxide (CO2), carbon monoxide (CO) and methanol (CH3OH), are locked in ices (boogert11; boogert13; oberg11) on the mantles of interstellar dust grains, in amounts predicted in some cases to far exceed those in the gas phase (hollenbach09). Ices play a major role in forming planetesimals within disks, and are a key source of water and organic molecules for newly forming planets. Grain mantles are the sites of active surface chemistry (tielens82; cuppen09), and thus one might expect ice composition to depend on their environment. The limited number of existing ice spectra is insufficient to determine whether the ices in protoplanetary disks, including our early Solar System, are mainly inherited from their progenitor clouds (dhooghe17), or whether ice composition evolves over time (aikawa12).
SPHEREx addresses these questions by conducting a comprehensive survey of ices to better understand their abundances, compositions, and how they evolve during the early phases of star and planet formation. We selected a list of more than 9 million targets based on 2MASS and WISE 3.4 and 4.6 µm catalogs using three criteria: 1) a line-of-sight extinction ; 2) a WISE 4.6 m flux brighter than 15 AB mag to ensure a high signal-to-noise ratio in each SPHEREx spectral channel; and 3) targets that are separated from any neighboring source contributing more than 1 % of their flux by at least 7″. The SPHEREx ices target list (Ashby23) will greatly expand – by more than a factor of 10,000 – the database of ice absorption spectra previously obtained by ISO, AKARI, Spitzer and JWST, significantly advancing our understanding of how ices form and evolve as gas and dust transition from diffuse to dense interstellar clouds, to circumstellar envelopes, and ultimately to protoplanetary disks (Melnick2025; see Figure 4).
@@ 2.4 All-Sky Spectral Survey
All-sky surveys (e.g. IRAS, COBE, AKARI, WISE) play a major role in advancing the field of astronomy, producing versatile legacy archives that prove valuable for decades. SPHEREx extends this tradition by providing the first all-sky near-infrared spectral survey. SPHEREx’s rich spectral database supports numerous scientific investigations (dore16). These range from studies of main sequence, low-mass, and evolved stars, to spectral catalogs of galaxies, to surveys of solar system objects, to observations of clusters, to Galactic maps of line and PAH emission features. The all-sky survey complements recent and forthcoming astrophysics satellites (dore18), including JWST, TESS, Gaia, Euclid, eROSITA and Roman.
@@ 2.5 Legacy Catalogs
In addition to data products that support the 3 core science themes (see §2.1 - §2.3), the SPHEREx science team is curating 3 spectral catalogs that benefit from the team’s specialized knowledge of the instrument: asteroids and comets; planet-bearing and low-mass stars; and clusters of galaxies. The solar system objects catalog of asteroids and comets consists of objects with known, but changing positions. ivezic22 estimates that SPHEREx will deliver meaningful flux measurements for about 100,000 asteroids, of which 10,000 objects will have high-quality spectra. Because the spectral observations are distributed over time, care must be taken to account for the effects of rotation and changing distance. The wavelengths span from scattered sunlight to thermal emission and include characteristic features of olivine, pyroxene, hydroxyl, water ice, and organics, and contain information on thermal properties that complement broadband photometry with the upcoming NEO Surveyor satellite (mainzer23).
SPHEREx will provide spectra to supplement external photometry from Gaia, GALEX, 2MASS and WISE for over 600,000 main sequence stars surveyed for exoplanets by the transit missions Kepler, K2 and TESS. These stellar SEDs will help determine best-fit effective temperatures, surface gravities, and [Fe/H] ratios to percent or sub-percent precision. When combined with Gaia astrometry, we can derive bolometric luminosities, proper motions, stellar radii and extinctions. This combination of data significantly improves planetary radius estimates for many transiting planets, by reducing the error from uncertainty in the stellar radius, which can otherwise dominate, especially for cool stars (stevens18). SPHEREx will also compile a spectral atlas of 10,000 low-mass stars, from late M dwarfs and all accessible brown dwarfs down to the coolest Y dwarfs. Low-resolution near-infrared spectroscopy is crucial for studying these cool stellar objects, which can have rich spectra due to their complex atmospheres. The observations will use Gaia astrometry to account for significant proper motions.
@@ 3 Science to Requirements
SPHEREx is designed to fulfill the science requirements of its three primary scientific objectives. In short, the inflationary cosmology theme is driven by having sufficient high-accuracy and low-accuracy redshifts to constrain the non-Gaussianity parameter through the bispectrum (Heinrich24) and power spectrum, respectively. These in turn set requirements for spectral resolving power and point source sensitivity in bands 1-4 (see Table 2). The history of galaxy formation theme can carry out its science with the same resolving power and sensitivity, but must control noise and systematic errors on degree angular scales in order to accurately map extended cosmological structures. The interstellar ices science theme studies bright Galactic sources, but requires higher resolving power, with wavelength coverage out to 5 to spectrally separate ice absorption features. We describe the flow of engineering requirements from science in abbreviated form, but note that extensive simulations were carried out to justify these choices during the mission development.
@@ 3.1 Spectral Coverage and Resolving Power
The four short-wavelength bands were designed for the cosmological redshift survey, with resolving powers of (see Table 2). We carried out simulations (Stickley16) of SPHEREx observations on the COSMOS field, using realistic galaxy spectra and including the effects of spectral confusion from faint galaxies below SPHEREx’s sensitivity. These studies indicated that we could obtain redshifts with accuracy up to , required for the bispectrum analysis, using only continuum features. The most recent study of SPHEREx redshifts (Feder24) confirms our choice of resolving power is appropriate, and indicates we could enhance the number of high-accuracy redshifts by incorporating emission lines. The galaxy formation theme can be accomplished with the wavelength and spectral resolution choices for cosmology.
@@ 3.2 Point Source Sensitivity
The inflationary cosmology theme constrains inflationary non-Gaussianity through measurements of the power spectrum and bispectrum with a large catalog of galaxies with a range of redshift accuracy (Dore14). The number density of galaxies as a function of redshift, their clustering biases, and their redshift accuracy set the measurement error on . The forecasting methodology used for the power spectrum is described in Dore14 and deputter17. The forecasting methodology used for the bispectrum follows Heinrich24 and includes an accurate handling of redshift errors, but excludes the non-Gaussian contribution to the covariance matrix (Biagetti:2021tua; Salvalaggio:2024) and the covariance between the power spectrum and bispectrum (dePutter:2018). It also assumes the galaxy biases universality relation (Dalal:2007cu; Barreira:2020ekm). All these effects, when included, might reduce the sensitivity by tens of percent. However, a full analysis incorporating a multi-tracer analysis with optimal samples may substantially improve the measurement error on from our initial forecast.
Feder24 studied the number and accuracy of redshifts using a template fitting method, which strongly depends on point source sensitivity in bands 1-4. In order to achieve the accuracy needed to meaningfully discriminate between single-field and multi-field inflation models, we set a requirement that the point source sensitivity exceed 18.4 AB mag () per spectral channel (see §4.2) in bands 1-4 once observations are complete (see Figure 6).